OpenStudy (anonymous):

What is the area of a sector with measure of arc equal to 90° and radius equal to 1 foot?

hero (hero):

Hi Cherri

OpenStudy (anonymous):

hi

hero (hero):

I don't know why you're making strange assumptions. No one is avoiding you

hero (hero):

anyways, I'm going to help you with this question

OpenStudy (anonymous):

no hero u misunderstood me i said were ur here wen im not and im hre wen ur not i didnt say anything abt adoviding

hero (hero):

\[\frac{x}{{\pi}r^2} = \frac{90}{360}\]

hero (hero):

Oh, I see. Well, now we're both here at the same time.

hero (hero):

\[\frac{x}{{\pi}(1)^2}=\frac{90}{360}\]

hero (hero):

\[90\pi = 360x \] \[x = \frac{90{\pi}}{360}\]

hero (hero):

\[x = \frac {\pi}{4}\]

hero (hero):

So the area of a sector is pi/4 units squared